OpenStudy (anonymous):

Determine whether the proportion is true or false. 15/6 = 5/2 is it true because 5 goes into 5 3x's & 2 goes into 6 3x's to?

OpenStudy (bahrom7893):

15/6 = 5*3/(2*3) = 5/2 yes that's true

OpenStudy (anonymous):

yes